OpenStudy (anonymous):

Find the exact value of the following limit: lim[x -> 0] (e*6x - 6x - 1)/(x^2)

hartnn (hartnn):

L'Hopital's twice

hartnn (hartnn):

bdw, whats the numerator exactly?

ganeshie8 (ganeshie8):

\[\lim\limits_{x\to 0}\frac{e^{6x} - 6x-1}{x^2}\] like this ?

OpenStudy (anonymous):

But when i do the rule one i get -6/0 which doesn't meet the conditions

OpenStudy (anonymous):

Yes, ganeshie8

hartnn (hartnn):

hmm (6e^(6x)-6) /2x it is 0/0 how are u getting -6 ?

hartnn (hartnn):

e^0=1

OpenStudy (anonymous):

but when you take the derivative doesn't the 6x turn into 6

OpenStudy (anonymous):

wait never mind i see where i made my mistake

hartnn (hartnn):

sure \(\Large (e^{6x})' = 6e^{6x} \\ \large (-6x)' =-6\)

hartnn (hartnn):

any more doubts? :)

OpenStudy (anonymous):

No more doubts i know how to do it now

hartnn (hartnn):

^_^

OpenStudy (anonymous):

Why don't you guys try it out the hard way. It's actually better learning

ganeshie8 (ganeshie8):

whats the hard way ?

OpenStudy (ikram002p):

the definition is bretty much easy :P

OpenStudy (anonymous):

Without Hopital's rule. Analytical solving..

ganeshie8 (ganeshie8):

epsilon delta ? for that you need to know the limit upfront right ?

ganeshie8 (ganeshie8):

ohk..

hartnn (hartnn):

the numerator ain't factorable

OpenStudy (anonymous):

Thank you guys so much for your help. XD

ganeshie8 (ganeshie8):

taylor series is also a kindof lhopital, so you dont want to use it either ?

ganeshie8 (ganeshie8):

this may be useful\[\large \lim\limits_{x\to 0}\dfrac{e^x-1}{x} = 1\]

OpenStudy (anonymous):

exactly

ganeshie8 (ganeshie8):

the denominator is x^2 however, so im not so sure

OpenStudy (freckles):

This is a really odd way... but \[u=x^2 \text{ as } x->0, u->0 \\ \pm \sqrt{u}=x \text{ so we have two cases to work } \\ \text{ using } x=\sqrt{u} \text{ we have } \\ f'(x^+)=f'(\sqrt{u})=\lim_{u \rightarrow 0} \frac{f(u)-f(0)}{u-0} \\ f(u)=e^{6 \sqrt{u}}-6 \sqrt{u} \\ f'(u)=\frac{6}{2 \sqrt{u}}e^{6 \sqrt{u}}-\frac{6}{2 \sqrt{u}} =\frac{3}{\sqrt{u}}e^{6 \sqrt{u}}-\frac{3}{\sqrt{u}}=\frac{3e^{6 \sqrt{u}}-3}{\sqrt{u}}=3 \frac{e^{6 x}-1}{x} \\ \text{ let } v=6x \\ \lim_{x \rightarrow 0^+} 3 \frac{e^{6x}-1}{x} = \lim_{v \rightarrow 0^+} 3 \frac{e^{v}-1}{\frac{v}{6}}=18(1)=18\] \[f'(x^-)=f'(- \sqrt{u}) \\ f(u)=e^{-6 \sqrt{u}}+6 \sqrt{u} \\ f'(u)=\frac{-6}{2 \sqrt{u}}e^{-6 \sqrt{u}} +\frac{6 }{2 \sqrt{u}} =-3\frac{ e^{ -6\sqrt{u}}-1}{\sqrt{u}}=3 \frac{e^{6x}-1}{x}\] so we still get the same limit 18 This is a little odd but it seems to work

hartnn (hartnn):

thank goodness there is L'Hopital :O

OpenStudy (freckles):

I don't know if that is the way the others were talking of or not

OpenStudy (ikram002p):

nice :)